# Bayesian evidence for two peaks in the sound speed in cold dense QCD

### v3

@@ Acknowledgment
In this appendix I prove eqs. 10 and 12 guarantee a peak in $C_{s}$ above NS densities. I begin with the case when $C_{s,\mathrm{mean}}^{\mathrm{UD}}<1/3$. The aim is to show eq. 12 ensures that the peak bounded by the constant $C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3$ segment is preceded by a minimum less than $C_{s,\mathrm{mean}}^{\mathrm{UD}}<1/3$, thus necessitates a second peak in $C_{s}$. The right hand side of eq. 12 is given by the maximally soft EOS with $C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3$ depicted as the solid red curve in fig. 3. If eq. 12 is to hold, the EOS must have segments lie above the solid red curve, i.e., pass through the shaded region. I then pick an arbitrary point along such an EOS in the shaded region, for instance point A in fig. 3, and label the density range preceding it as region I and the one subsequent to it as region II. The mean value EOSs between point A and the TOV point and between point A and the pQCD point are shown as the solid black lines in fig. 3, and their associated mean values are labeled as $C_{s,\mathrm{mean}}^{\mathrm{I}}$, $C_{s,\mathrm{mean}}^{\mathrm{II}}$. The mean value bound eq. 6 informs the respective global minima and global maxima within regions I and II, requiring
@@
$C_{s,\mathrm{min}}^{\text{II}}$ $\leq C_{s,\mathrm{mean}}^{\text{II}}\leq C_{s,\mathrm{max}}^{\text{II}}.$ (16)
Among the two maxima (minima), at least one of them is the global maximum (minimum) across the entire ultra-dense phase $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ ($C_{s,\mathrm{min}}^{\mathrm{UD*}}$).
Next, I show that $C_{s,\mathrm{mean}}^{\text{I}}$ and $C_{s,\mathrm{mean}}^{\text{II}}$ are themselves constrained. These bounds can be obtained by comparing the slopes of these EOSs in fig. 3 to known values and recalling that the inverse of the slopes are the speed of sound squared (see eq. 4). Since the mean value EOS in region I is steeper than the mean value EOS over the entire range (underlying $C_{s,\mathrm{mean}}^{\mathrm{UD}}$), and the mean value EOS in region II is shallower than the red segment with constant $C_{s}=C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3$, one finds
@@
i.e., eqs. 1 and 3.
@@
Upon substituting in eq. 15 one recovers eq. 19. Even though this scenario where $C_{s,\mathrm{mean}}^{\mathrm{UD}}>1/3$ is disfavored by the existence of two-solar-mass NSs, it is still included because (i) in the Bayesian framework eventually excluded samples still require a physical likelihood to be evaluated; and (ii) it is relevant if ${\chi\text{EFT}}$ breaks down earlier, for instance due to a low-density quarkyonic transition. Analyses that follow ${\chi\text{EFT}}$ to lower densities are presented in appendix C.
@@ B.3 Interpolation Methods
  • Splines (linear, quadratic, cubic, B-spline, $\ldots$) are perhaps the simplest choices. My approach differs from those available in the market in that I do not impose the number of pieces, whose physical meaning is obscure. Instead, I directly work with the length scale $l\equiv\Delta x$ of the underlying thermodynamic variable $x$. $l(x)$ is generally density-dependent.
@@
    The strategy for choosing the length scales $l_{a}$ is similar to that in the previous section except here $l_{a}$ is independent of $x$. The strength of each Gaussian kernel $\sigma(l)$ is either taken to be completely random or follow a power law $\sigma(l)\propto l^{\beta}$ where the exponent $\beta$ is sampled from a flat distribution on $[-1,1]$. A negative (positive) exponent stresses shorter (longer) correlation lengths, and $\beta=0$ reflects a flat spectrum. I normalize the strengths by randomly picking $\Sigma=\sum_{a}\sigma(l_{a})$ from a uniform distribution on $[\Sigma_{\mathrm{min}},\Sigma_{\mathrm{max}}]$. Larger $\Sigma$ leads to more volatile features whereas lower $\Sigma$ tends to produce flatter sound speed. $\Sigma_{\mathrm{min}}$ and $\Sigma_{\mathrm{max}}$ are chosen to ensure a wide range of variances in $C_{s}$ are realized. The mean $\bar{y}$ is sampled randomly from a standard normal distribution $\bar{y}\sim\mathcal{N}(0,1)$. Finally, a white noise (not shown in eq. 31) of strength $\sigma=10^{-4}$ is added to the kernel for numerical stability.
@@ B.5 Summary of models
This appendix examines assumptions underlie the main results in addition to those discussed in appendix B. I begin by discussing choices of prior distributions. In the main text, neutral priors without preferences for the criterion $\Delta P-\Delta P_{\mathrm{mean,max}}$ are chosen. They are obtained by imposing separate flat distributions over regions where $\Delta P-\Delta P_{\mathrm{mean,max}}>0$ and where $\Delta P-\Delta P_{\mathrm{mean,max}}<0$. Normalizations of the two segments are obtained from demanding the integrated probability equals on both sides. The effect of eq. 7 is minor, and I have verified that neutral priors on $\Delta P-\Delta P_{\mathrm{mean,max}}$ alone lead to unbiased priors for the two-peak scenario.
@@
It is noteworthy that the criteria for the two-peak scenario eq. 13 are not equivalent to a single requirement on the maximum sound speed $C_{s,\mathrm{max}}^{\mathrm{UD*}}>1/3$. For instance, a monotonically decreasing $C_{s}^{\mathrm{UD}}$ with $C_{s,\mathrm{max}}^{\mathrm{UD*}}>1/3$ does not lead to a second peak at high densities. In fact, monotonically decreasing $C_{s}$ between TOV and pQCD points are strongly disfavored, see table 7 in appendix D. The upshot is that flat priors on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ are not the least informative ones. Nevertheless, it is illuminating to explicitly exclude high sound speed by imposing upper bounds on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ in the prior. For the purpose of clarifying phase space volumes, let us demand $C_{s,\mathrm{max}}^{\mathrm{UD*}}\leq 2/3$ by replacing the requirement on the prior eq. 14 with the following:
@@
The expression for $\Delta P_{\mathrm{min}}$ is related to eq. 11 via swapping subscripts TOV and pQCD and is given in ref [66]. The resulting Bayes factors are shown in table 4. Narrowing the phase-space volume in the prior reduces the evidence, but the main findings remain unaffected.
A side note on the treatment of eq. 14. In Bayesian analyses in the literature, e.g. refs [62, 63, 34], eq. 14 is often treated as a likelihood. Since eq. 14 simply comes from thermodynamic stability and causality, a priori one expects it to hold. Therefore, I consider it as a requirement on the prior, the same way causality and stability are built into most parameterizations of NS EOSs. Furthermore, because fully model-independent studies in the pressure-energy density [61] (albeit only a sufficient condition) and the mass-radius plane [64] are available, the effects of eq. 14 are robustly reflected as the boundaries that delimit physical regions in the phase spaces. Treating eq. 14 as a prior implicitly incorporates these clean and simple results. As such, the choice made here appears more sensible at least for the present purpose, both from a physical viewpoint and a Bayesian perspective.
@@
The phase space for the two-peak scenario is also controlled by the lower bound of sound speed squared $C_{s,\mathrm{min}}^{\mathrm{UD*}}$. So far, only $C_{s,\mathrm{min}}^{\mathrm{UD*}}=0$ required by thermodynamic stability is imposed. Low sound speed generally signals first-order phase transitions where low- and high-density phases in mechanical equilibrium are separated by a finite latent heat so that $C_{s}=\mathrm{d}P/\mathrm{d}\mathcal{E}\simeq 0$. Based on similar symmetry breaking patterns and low-lying excitations in nucleon superfluid and in quark color superconductor, it has been conjectured that the quark-hadron transition at finite density is a smooth crossover [87, 88, 89] 555though topological phase transitions are possible [114, 115, 116, 117]. The ultra-dense regime would still be free of phase transitions if they occur inside NSs., motivating more restrictive $C_{s,\mathrm{min}}^{\mathrm{UD*}}$. As discussed in the main text, when $C_{s,\mathrm{min}}^{\mathrm{UD*}}>0$ the area $\Delta P_{\mathrm{max}}(C_{s,\mathrm{min}}^{\mathrm{UD*}},C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3)$ decreases rendering the criterion eq. 12 less demanding. The Bayes factors as functions of $C_{s,\mathrm{min}}^{\mathrm{UD*}}$ are shown in figs. 7 and 8. With $C_{s,\mathrm{min}}^{\mathrm{UD*}}\gtrsim 0.1-0.15$, even the unfavorable data $M_{\text{TOV}}\lesssim 2.2M_{\odot}$ would yield substantial support for two peaks. When $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, most EOSs are incompatible with $C_{s,\mathrm{min}}^{\mathrm{UD*}}\gtrsim 0.1-0.2$ [66], rendering the statistics less reliable with stronger assumptions about $C_{s,\mathrm{min}}^{\mathrm{UD*}}$.
@@
Other parameters of this framework are $n_{\chi\text{EFT}}$, the highest density ${\chi\text{EFT}}$ is adopted, and $\mu_{\text{pQCD}}$, the lowest chemical potential pQCD remains reliable. The convergence of ${\chi\text{EFT}}$ up to $n_{\chi\text{EFT}}=2n_{0}$ is demonstrated by statistical analyses exploiting the structure of EFT [85], although its uncertainty may have been underestimated [118]. The evidence is only marginally affected if the EFT error band is broadened from $2\sigma$ to $4\sigma$, see table 6. Here and throughout this work, flat distributions within the ${\chi\text{EFT}}$ uncertainty band are assumed. This is achieved by imposing flat priors on the coefficients of basis functions in our ${\chi\text{EFT}}$-based low-density model, see eq 2 in [64]. Additionally, ${\chi\text{EFT}}$ may not be suitable if the ground state is no longer nucleonic at $\sim 2n_{0}$, such as the case of low-density quarkyonic transitions [40]. To examine these possibilities, I repeat the Bayesian analyses but only adopt ${\chi\text{EFT}}$ up to $n_{\chi\text{EFT}}=n_{0},1.5n_{0}$. The updated Bayes evidence is shown in table 6 and fig. 7. If the NS maximum mass is indeed low near $2.2M_{\odot}$, low-density nuclear theory inputs between $n_{0}$ and $2n_{0}$ would be helpful to secure clear evidence for two peaks in $C_{s}$.
@@
On the high-density end, $\mu_{\text{pQCD}}=2.4-2.6$ GeV is commonly adopted in the literature, although the large sensitivity on $X$ raises questions about this assumption. Refs  [55, 59] resummed to all order the leading and next-to-leading soft modes, and found reduced dependence on $X$ favoring $X\gtrsim 1.5$. Ref [34] inferred $X$ from comparing the N2LO pQCD EOS with lattice calculations at finite isospin [86, 33]. By including the LO pQCD prediction for the superconducting gap and exploiting the similar structure of pQCD at finite isospin and at finite $\mu_{B}$, $X$ is found to be large $X\in[2.4,4]$ and the superconducting gap at finite $\mu_{B}$ is found to be negligible with $\Delta\lesssim\text{MeV}$ near $\mu_{B}=2.4~{}\text{GeV}$. This bound on $X$ translates to a tightly constrained $P_{\text{pQCD}}\in[2.8,3.0]$ GeV/fm3 at $\mu_{B}=2.4~{}\text{GeV}$, which roughly corresponds to $X\gtrsim 2.3$ when soft modes at N3LO are included (the pQCD EOS adopted in the main text). According to table 1, such high values of $X$ lead to a notable preference for the two-peak scenario even with $M_{\text{TOV}}\lesssim 2.2M_{\odot}$. On the other hand, the suppressed superconducting gap seems to suggest other mechanisms are behind the postulated second peak in $C_{s}$. If the superconducting gap is underestimated by LO pQCD, the inferred $X$ thus $P_{\text{pQCD}}$ is reduced when matching to lattice data at finite isopsin, but the full pressure $P_{\mathrm{CQM}}=P_{\text{pQCD}}+P_{\Delta}$ receives a pairing contribution $P_{\Delta}=\mu_{B}^{2}\Delta^{2}/(3\pi^{2})$ that is important if $\Delta\gtrsim 100$ MeV and $X\lesssim 1.5$. The bottom line is that pairing gaps much larger than inferred in ref [34] can still yield a large $\Delta P$ essential for the two-peak evidence, even though $X$ would be reduced. These large gaps are near the high end of current estimates, and are compatible with astrophysics [64, 65]. In short, while there exists strong indications supporting the reliability of pQCD down to $\mu_{B}=2.4$ GeV, higher order calculations and detailed statistical analyses will be essential to assess the applicable range of asymptotic expansions in $\alpha_{s}$ and to pinpoint the size of superconducting gaps.

### v4

@@ Acknowledgment
In this appendix I prove eqs. 10 and 12 guarantee a peak in $C_{s}$ above NS densities, and that along with eq. 7 they ensure two peaks at supranuclear densities. I begin with the case when $C_{s,\mathrm{mean}}^{\mathrm{UD}}<1/3$. The aim is to show eq. 12 ensures that the peak bounded by the constant $C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3$ segment is preceded by a minimum less than $C_{s,\mathrm{mean}}^{\mathrm{UD}}$. The right hand side of eq. 12 is given by the maximally soft EOS with $C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3$ depicted as the solid red curve in fig. 3. If eq. 12 holds, the EOS must have segments lie above the solid red curve, i.e., pass through the shaded region. I then pick an arbitrary point along such an EOS in the shaded region, for instance point A in fig. 3, and label the density range preceding it as region I and the one subsequent to it as region II. The mean value EOSs between point A and the TOV point and between point A and the pQCD point are shown as solid black lines in fig. 3, and their associated mean values are labeled as $C_{s,\mathrm{mean}}^{\mathrm{I}}$, $C_{s,\mathrm{mean}}^{\mathrm{II}}$. The mean value bound eq. 6 informs the respective global minima and global maxima within regions I and II, requiring
@@
$C_{s,\mathrm{min}}^{\text{II}}$ $\leq C_{s,\mathrm{mean}}^{\text{II}}\leq C_{s,\mathrm{max}}^{\text{II}}.$
Next, I show that $C_{s,\mathrm{mean}}^{\text{I}}$ and $C_{s,\mathrm{mean}}^{\text{II}}$ are themselves constrained. These bounds can be obtained by comparing the slopes of these EOSs in fig. 3 to known values and recalling that the inverse of slopes are the speed of sound squared (eq. 4). Since the mean value EOS in region I is steeper than the mean value EOS over the entire ultra-dense phase (underlying $C_{s,\mathrm{mean}}^{\mathrm{UD}}$), and the mean value EOS in region II is shallower than the red segment with constant $C_{s}=C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3$, one finds
@@
i.e., eqs. 1 and 3. And since the minimum in region I is below $C_{s,\mathrm{mean}}^{\mathrm{UD}}$, eq. 7 implies eq. 2 regardless of the number of minima in the ultra-dense phase.
@@
Upon substituting in eq. 15 one again recovers
$C_{s,\mathrm{min}}^{\text{I}}$ $<C_{s,\mathrm{mean}}^{\mathrm{UD}},$ (22)
$C_{s,\mathrm{max}}^{\text{II}}$ $>1/3.$ (23)
Even though this scenario where $C_{s,\mathrm{mean}}^{\mathrm{UD}}>1/3$ is disfavored by the existence of two-solar-mass NSs, it is still included because (i) in the Bayesian framework eventually excluded samples still require a physical likelihood to be evaluated; and (ii) it is relevant if ${\chi\text{EFT}}$ breaks down earlier, for instance due to a low-density quarkyonic transition [41]. Analyses that follow ${\chi\text{EFT}}$ to lower densities are presented in appendix C.
@@ B.3 Interpolation Methods
  • Splines (linear, quadratic, cubic) are perhaps the simplest choices. My approach differs from those available in the market in that I do not impose the number of pieces, whose physical meaning is obscure. Instead, I directly work with the length scale $l\equiv\Delta x$ of the underlying thermodynamic variable $x$. $l(x)$ is generally density-dependent.
@@
    The strategy for choosing the length scales $l_{a}$ is similar to that in the previous section except here $l_{a}$ is independent of $x$. I fix the grid spacing $\Delta x$ to be the lesser of $\min\{l_{a}\}/3$ and $(x_{\mathrm{max}}-x_{\mathrm{{\chi\text{EFT}}}})/500$, so that the resolution is adequate for short correlation lengths. The strength of each Gaussian kernel $\sigma(l)$ is either taken to be completely random or follow a power law $\sigma(l)\propto l^{\beta}$ where the exponent $\beta$ is sampled from a flat distribution on $[-1,1]$. A negative (positive) exponent stresses shorter (longer) correlation lengths, and $\beta=0$ reflects a flat spectrum. I normalize the strengths by randomly picking $\Sigma=\sum_{a}\sigma(l_{a})$ from a uniform distribution on $[\Sigma_{\mathrm{min}},\Sigma_{\mathrm{max}}]$. Larger $\Sigma$ leads to more volatile features whereas lower $\Sigma$ tends to produce flatter sound speed. $\Sigma_{\mathrm{min}}$ and $\Sigma_{\mathrm{max}}$ are chosen to ensure a wide range of variances in $C_{s}$ are realized. The mean $\bar{y}$ is sampled randomly from a standard normal distribution $\bar{y}\sim\mathcal{N}(0,1)$. Finally, a white noise (not shown in eq. 32) of strength $\sigma=10^{-4}$ is added to the kernel for numerical stability.
@@ B.5 Summary of models
This appendix examines assumptions underlie the main results in addition to those discussed in appendix B. I begin by discussing choices of prior distributions. In the main text, neutral priors without preferences for the criterion $\Delta P-\Delta P_{\mathrm{mean,max}}$ are chosen. They are obtained by imposing separate flat distributions over regions where $\Delta P-\Delta P_{\mathrm{mean,max}}>0$ and where $\Delta P-\Delta P_{\mathrm{mean,max}}<0$. The two segments are normalized such that their integrated probability equals. The effect of eq. 7 is minor, and I have verified that neutral priors on $\Delta P-\Delta P_{\mathrm{mean,max}}$ alone lead to unbiased priors for the two-peak scenario.
@@
It is noteworthy that the criteria for the two-peak scenario eq. 13 are not equivalent to a single requirement on the maximum sound speed $C_{s,\mathrm{max}}^{\mathrm{UD*}}>1/3$. For instance, a monotonically decreasing $C_{s}^{\mathrm{UD}}$ with $C_{s,\mathrm{max}}^{\mathrm{UD*}}>1/3$ does not lead to a second peak at high densities. In fact, monotonically decreasing $C_{s}$ between TOV and pQCD points are strongly disfavored, see table 7 in appendix D. The upshot is that flat priors on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ are not the least informative ones. Nevertheless, it is illuminating to explicitly exclude high sound speed by imposing upper bounds on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ in the prior. For instance, let us demand $C_{s,\mathrm{max}}^{\mathrm{UD*}}\leq 2/3$ by replacing the requirement on the prior eq. 14 with the following:
@@
The expression for $\Delta P_{\mathrm{min}}$ is related to eq. 11 via swapping $\alpha$ and $\beta$ and is given in ref [67]. The resulting Bayes factors are shown in table 4. Narrowing the phase-space volume in the prior reduces the evidence, but the main findings remain unaffected.
A side note on the treatment of eq. 14. In Bayesian analyses in the literature, e.g. refs [63, 64, 65, 35], eq. 14 is often treated as a likelihood. Since eq. 14 simply comes from thermodynamic stability and causality, a priori one expects it to hold. Therefore, I consider it a requirement on the prior, the same way causality and stability are built into most parameterizations of NS EOSs. Furthermore, because fully model-independent studies of the pressure-energy density relation [62] and in the mass-radius plane [65] are available, the effects of eq. 14 are robustly determined as the boundaries that delimit physical regions in the phase spaces. Treating eq. 14 as a prior implicitly incorporates these clean and simple results. As such, the choice made here appears to be more sensible, at least for the present purpose, both from a physical viewpoint and a Bayesian perspective.
When taken as a likelihood, eq. 14 is found to favor soft EOSs in NS cores [63, 64, 35] since acausal extensions in the ultra-dense phase are more common among stiff NS cores. When treated as a prior, almost any arbitrarily stiff core can be compatible with eq. 14 [65]. To sum up, stiffening of the EOS in the density range $\sim 2-5n_{0}$ is a physical consequence of two-solar-mass pulsars [4, 40], and only additional constraints on $M_{\text{TOV}}$ can robustly inform the NS EOS at higher densities.
@@
The phase space for the two-peak scenario is also controlled by the lower bound of sound speed squared $C_{s,\mathrm{min}}^{\mathrm{UD*}}$. So far, only $C_{s,\mathrm{min}}^{\mathrm{UD*}}=0$ required by thermodynamic stability is imposed. Low sound speed generally signals first-order phase transitions where low- and high-density phases in mechanical equilibrium are separated by a finite latent heat so that $C_{s}=\mathrm{d}P/\mathrm{d}\mathcal{E}\simeq 0$. Based on similar symmetry breaking patterns and low-lying excitations in nucleon superfluid and in quark color superconductor, it has been conjectured that the quark-hadron transition at finite density is a smooth crossover [88, 89, 90] 666though topological phase transitions are possible [115, 116, 117, 118]. The ultra-dense regime would still be free of phase transitions if they occur inside NSs., motivating more restrictive $C_{s,\mathrm{min}}^{\mathrm{UD*}}$. As discussed in the main text, when $C_{s,\mathrm{min}}^{\mathrm{UD*}}>0$ the area $\Delta P_{\mathrm{max}}(C_{s,\mathrm{min}}^{\mathrm{UD*}},C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3)$ decreases rendering the criterion eq. 12 less demanding. The Bayes factors as functions of $C_{s,\mathrm{min}}^{\mathrm{UD*}}$ are shown in figs. 7 and 8. With $C_{s,\mathrm{min}}^{\mathrm{UD*}}\gtrsim 0.1-0.15$, even the unfavorable data $M_{\text{TOV}}\lesssim 2.2M_{\odot}$ would yield substantial support for two peaks. When $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, most EOSs are incompatible with $C_{s,\mathrm{min}}^{\mathrm{UD*}}\gtrsim 0.1-0.2$ [67], rendering the statistics less reliable with stronger assumptions about $C_{s,\mathrm{min}}^{\mathrm{UD*}}$.
@@
Other parameters of this framework are $n_{\chi\text{EFT}}$, the highest density ${\chi\text{EFT}}$ is adopted, and $\mu_{\text{pQCD}}$, the lowest chemical potential pQCD remains reliable. The convergence of ${\chi\text{EFT}}$ up to $n_{\chi\text{EFT}}=2n_{0}$ is demonstrated by statistical analyses exploiting the structure of EFT [86], although its uncertainty may have been underestimated [119]. The evidence is only marginally affected if the EFT error band is broadened from $2\sigma$ to $4\sigma$, see table 6. Here and throughout this work, flat priors within the ${\chi\text{EFT}}$ uncertainty band are imposed by sampling uniformly the coefficients of basis functions in our ${\chi\text{EFT}}$-based low-density model, see eq 2 in [65]. Additionally, ${\chi\text{EFT}}$ may not be suitable if the ground state is no longer nucleonic at $\sim 2n_{0}$, such as the case of low-density quarkyonic transitions [41]. To examine these possibilities, I repeat the Bayesian analyses but only adopt ${\chi\text{EFT}}$ up to $n_{\chi\text{EFT}}=n_{0},1.5n_{0}$. The updated Bayes evidence is shown in table 6 and fig. 7. If the NS maximum mass is indeed low near $2.2M_{\odot}$, low-density nuclear theory inputs between $n_{0}$ and $2n_{0}$ would be helpful to secure clear evidence for two peaks in $C_{s}$.
@@
On the high-density end, $\mu_{\text{pQCD}}=2.4-2.6$ GeV is commonly adopted in the literature, although the large sensitivity on $X$ raises questions about this assumption. Refs  [56, 60] resummed to all order the leading and next-to-leading soft modes, and found reduced dependence on $X$ favoring $X\gtrsim 1.5$. Ref [35] inferred $X$ from comparing the N2LO pQCD EOS with lattice calculations at finite isospin [87, 34]. By including the LO pQCD prediction for the superconducting gap and exploiting the similar structure of pQCD at finite isospin and at finite $\mu_{B}$, $X$ is found to be large $X\in[2.4,4]$ and the superconducting gap at finite $\mu_{B}$ is found to be negligible with $\Delta\lesssim\text{MeV}$ near $\mu_{B}=2.4~{}\text{GeV}$. This bound on $X$ translates to a tightly constrained $P_{\text{pQCD}}\in[2.8,3.0]$ GeV/fm3 at $\mu_{B}=2.4~{}\text{GeV}$, which roughly corresponds to $X\gtrsim 2.3$ when soft modes at N3LO are included (the pQCD EOS adopted in the main text). According to table 1, such high values of $X$ lead to a notable preference for the two-peak scenario even with $M_{\text{TOV}}\lesssim 2.2M_{\odot}$. On the other hand, the suppressed superconducting gap seems to suggest that other mechanisms are behind the postulated second peak in $C_{s}$. If the superconducting gap is underestimated by LO pQCD, the inferred $X$ thus $P_{\text{pQCD}}$ is reduced when matching to lattice data at finite isopsin, but the full pressure $P_{\mathrm{CQM}}=P_{\text{pQCD}}+P_{\Delta}$ receives a pairing contribution $P_{\Delta}=\mu_{B}^{2}\Delta^{2}/(3\pi^{2})$ that is important if $\Delta\gtrsim 100$ MeV and $X\lesssim 1.5$. The bottom line is that pairing gaps much larger than inferred in ref [35] can still yield a large $\Delta P$ essential for the two-peak evidence, even though inferred $X$ would be reduced. These large gaps are near the high end of current estimates, and are compatible with astrophysics [65, 66]. In short, while there are indications that pQCD is reliable down to $\mu_{B}=2.4$ GeV, higher-order calculations and detailed statistical analyses will be essential to confirm the applicable range of asymptotic expansions in $\alpha_{s}$ and to pinpoint the size of superconducting gaps.